# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ 7 Numerical experiments
This section presents some numerical experiments to compare the accuracy and run times of Algorithm 1 with some state-of-the-art algorithms for computing regularization and solution paths to (BPDN). Specifically, we compare below our algorithms with the glmnet software package (glmnet[61, 76, 30, 37], the MATLAB’s native lasso implementation [^3] This is ostensibly an implementation of the glmnet package to MATLAB, except that it has more options, including for better control of the accuracy. (mlasso), and the fast iterative shrinkage thresholding algorithm (fista) with a strong selection rule [6, 32, 62, 53]. For the special case $t=0$, we compare the performance of Algorithm 1 with regularization path starting from $(t^{(0)},\boldsymbol{p}^{(0)})=(\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty},-\boldsymbol{b}/t^{(0)})$, no regularization path and input $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, and no regularization path and input provided by Algorithm 3. All experiments below were carried out in MATLAB on an Apple M3 processor.
@@
All our algorithms require solving a NNLS problem. For this, we use Meyer’s algorithm [47], an active set NNLS algorithm generalizing the Lawson–Hanson algorithm [41] to arbitrary starting active sets, and so particularly suitable for BPDN. For glmnet, we use the code available at https://github.com/junyangq/glmnet-matlab. For fista, we use the variant described in [16, Algorithm 5, Page 197] together with the selection rule described in [62, Equation 7]. For our numerical experiments, we use the benchmark datasets provided by Lorenz et al. [44, Tables 1-2, Section 4], tailored specifically for BPDN and BP and to use for comparing different numerical algorithms. Each dataset comprises a triplet $(\bm{A},\boldsymbol{b},\boldsymbol{x}_{\text{BP}}^{s})$, where $\boldsymbol{x}_{\text{BP}}^{s}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{m}}\left\|{\boldsymbol{x}}\right\|_{1}\ \text{s.t.}\ \bm{A}\boldsymbol{x}=\boldsymbol{b}$. We use in total ten different datasets, corresponding to some of their largest dense matrix ($m=1024,\,n=8192$) with six different observed data and solution vectors and their largest sparse matrix ($m=8192,\,n=49152$) with four different observed data and solution vectors. The solution vectors have either high dynamic range (HDR) or low dynamic range (LDR), and their support either satisfy the so-called Exact Recovery Condition (ERC) [65], an extended form of it (extERC) or, for dense matrices only, no exact recovery condition (noERC) (see [44] for details). All our MATLAB implementations, external software and datasets will be bundled and made publicly available on Zotero.
@@ 7.1 Accuracy checks
We compare how accurately glmnet, fista and mlasso compute solutions to (BPDN) and (dBPDN) (via the optimality conditions $t\boldsymbol{p}^{s}(t,\boldsymbol{b})=\bm{A}\boldsymbol{x}^{s}(t,\boldsymbol{b})-\boldsymbol{b}$) across solution paths. We use dataset 548 (dense $1024\times 8192$ matrix, LDR, noERC) and dataset 474 (sparse $8192\times 49512$ matrix, LDR, extERC) from [44], their most computationally demanding datasets. For each dataset, we first compute the entire solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ using Algorithm 2 and identify all $K$ kinks in $t$ in the solution path (see Remark 5.1). Then we run Algorithm 1, glmnet, fista, and mlasso (for dataset 584 as mlasso does not support sparse matrices) at the kinks $\{t^{(k)}\}_{k=0}^{K}$, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$.
We perform two runs for each dataset. In the first run, we use the default options of glmnet and mlasso and a relative tolerance of $10^{-4}$ for fista (i.e., the algorithm stops when the relative difference in $\ell_{\infty}$-norm of the dual updates is less than $10^{-4}$). In the second run, we use harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista). In both runs, we use a tolerance of $10^{-8}$ in Algorithms 1 and 2 when evaluating their equicorrelation sets on line 3. After convergence of each algorithm, we evaluate the dual objective function in (dBPDN) and number of nonzero components of their primal solutions along the solution path.
@@
At default tolerance, mlasso and fista performed reasonably well while glmnet performed poorly. In particular, glmnet produced a numerical solution to BPDN with many more nonzero components than all other algorithms. This even remained the case in dataset # 474 with a harsh tolerance: glmnet produced a numerical solution to BPDN with a higher number of nonzero components compared to fista and Algorithm 1. Taken together, we see that Algorithm 1 is the clear winner when it comes to accuracy, with mlasso and fista performing well, and with glmnet performing poorly, even when using harsher tolerances.
Figure 1: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution paths. The dataset used is # 548 from [44]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-04}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
Figure 2: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution path. The dataset used is # 474 from [44]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-04}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
@@ 7.2 Run times comparisons
7.2.1 Computing regularization paths to basis pursuit denoising
Table 1 shows the run times of Algorithm 1, glmnet, mlasso, and fista for computing regularization paths. The regularization paths were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$. We also included the point $\{0\}$ for Algorithm 1, which the other algorithms cannot handle. Since we wish to compute accurate regularization paths, we used harsh tolerances for each algorithm, just as in Section 7.1 ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-08}$ for mlasso, and a relative tolerance of $081e-$ 08 for fista). Table 1 shows that the overall winner is Algorithm 1; it is significantly faster than fista and mlasso, and slightly faster or comparable to glmnet.
@@ 7.2.1 Computing regularization paths to basis pursuit denoising
7.2.2 Computing optimal solutions to basis pursuit
Table 2 shows the run times of Algorithm 1 with $t=0$ and using as input the feasible dual solution $\boldsymbol{p}_{f}$ obtained from Algorithm 3, Algorithm 1 with $t=0$ and using as input $-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, and Algorithm 1 with a regularization path. We used the ten aforementioned benchmark datasets from [44], and the regularization paths for Algorithm 1 were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$ and including $\{0\}$. We did not include MATLAB’s native linear programming solver because we found that it was slow and often failed to converge. We also did not include other solvers for (BP), such as AISS [14], SPGL1 [67] and SeDuMi [59], because earlier work by Tendero et al. [60] did similar comparisons with a different version of Algorithm 1 (valid for $t=0$ only) and found it superior in performance and accuracy.
Table 2 shows that all the methods have similar running times. In addition, we’ve verified that all methods correctly computed the solution vectors provided in the benchmark datasets. Algorithm 1 with regularization path is often slower because it used a large number of points. Using Algorithm 3 to get a warm input for Algorithm 1 performs just as well as using Algorithm 1, and in dataset # 274 it was slightly faster.
Datasets Alg. 1 (RP) Alg. 1 ($t=0$) Algs. 1 and 3
# 147 (d, HDR, ERC) 0.65 0.019 0.030
# 148 (d, HDR, extERC) 0.69 0.041 0.047
# 274 (d, HDR, noERC) 1.22 0.72 0.51
# 421 (d, LDR, ERC) 0.58 0.020 0.025
# 422 (d, LDR, extERC) 0.60 0.035 0.044
# 548 (d, LDR, noERC) 1.06 0.58 1.69
# 199 (s, HDR, ERC) 0.65 0.028 0.034
# 200 (s, HDR, extERC) 1.07 1.23 1.20
# 473 (s, LDR, ERC) 0.50 0.028 0.036
# 474 (s, LDR, extERC) 0.98 1.29 1.93
Table 2: Timings (in seconds) for Algorithm 1 with a regularization path, Algorithm 1 with $t=0$ and using $-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$ as input, and Algorithm 1 with $t=0$ and using the feasible dual solution $\boldsymbol{p}_{f}$ obtained from Algorithm 3 as input. Total times for 512/1024 logarithmically spaced points and including $\{0\}$ (d: dense/s: sparse), averaged over 5 runs.
@@ 8 Conclusion and future work
In this work, we proved that a minimal selection principle from the theory of differential inclusions (Theorem 3.1) enables one to compute an optimal solution of (dBPDN) from the asymptotic limit of the slow system (12). As the results in Section 3 and 4 show, the slow system can be integrated, yielding the slow solution (30). The slow solution converges in finite time and, at convergence, yields the optimal solution to (dBPDN). From it, one can recover an optimal solution to (BPDN). Taken together, these results yielded Algorithm 1. We also presented, in Section 5, a detailed perturbation analysis of the slow system, including its local and global dependence on the hyperparameter and data. The global continuation of the slow solution provided a rigorous, generalized homotopy algorithm for BPDN, as well as Algorithm 3, a greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time. Our numerical experiments showed that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting its key feature that it neither compromises accuracy nor computational efficiency.
While this work focused largely on BPDN, our results provided a novel solution method for solving a broad class of projected dynamical systems. We therefore expect that this method will be relevant to applications involving variational inequalities and projected dynamical systems. In addition, we expect that our results can be adapted to compute exact or approximate solutions to a broader class of convex polyhedral-constrained optimization problems. This includes: (i) variations of BPDN where the $\ell_{1}$ norm is replaced by a polyhedral norm or the function $\boldsymbol{x}\mapsto\left\|{\bm{M}\boldsymbol{x}}\right\|_{1}$ for some appropriate real matrix $\bm{M}$, (ii) other $\ell_{1}$-regularized problems such as logistic regression, Poisson regression, support vector machines and boosting problems, and (iii) extensions to inequality constraints. In particular, as basis pursuit and linear programming are theoretically equivalent problems [64], the results of this work suggest it may be possible to devise algorithms for linear programming based on differential inclusions. These problems will be investigated in future work.
@@ Appendix A Mathematical background
We list here important definitions and technical results from convex and functional analysis used in this work. For comprehensive references, we refer the reader to [54, 55, 3, 26, 38, 39].
Notation
All vectors and matrices are denoted in bold typeface. Given an $m\times n$ real matrix $\bm{A}$ and a set of indices $\mathcal{E}\subset\{1,\dots,n\}$, we write $\bm{A}_{\mathcal{E}}$ to denote its $m\times|\mathcal{E}|$ submatrix with columns indexed by $\mathcal{E}$ and we write $\bm{A}^{\top}_{\mathcal{E}}$ to denote its transpose. Similarly, given $\boldsymbol{u}\in\mathbb{R}^{n}$, we write $\boldsymbol{u}_{\mathcal{E}}$ to denote its $|\mathcal{E}|$-dimensional subvector indexed by $\mathcal{E}$.
@@
A nonempty set $C\subset\mathbb{R}^{n}$ is a closed convex polyhedron if it can be expressed as
where $\{\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{l}\}\subset\mathbb{R}^{n}$ and $\{r_{1},\dots,r_{l}\}\subset\mathbb{R}$.
@@
A function $f$ defined over $\mathbb{R}^{n}$ is proper if its domain
$\mathrm{dom~{}}f\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:f(\boldsymbol{x})<+\infty\right\}$
is nonempty and $f(\boldsymbol{x})>-\infty$ for every $\boldsymbol{x}\in\mathrm{dom~{}}f$.
@@
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is lower semicontinuous at $\boldsymbol{x}\in\mathrm{dom~{}}f$ if for every sequence $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ converging to $\boldsymbol{x}$,
$\liminf_{k\to+\infty}f(\boldsymbol{x}_{k})\geqslant f(\boldsymbol{x}).$
We say that $f$ is lower semicontinuous if it is lower semicontinuous at every $\boldsymbol{x}\in\mathrm{dom~{}}f$.
@@
Let $f$ be a proper function with $\mathrm{int}{~{}\mathrm{dom~{}}f}\neq\emptyset$. The function $f$ is differentiable at $\boldsymbol{x}\in\mathrm{int}{~{}\mathrm{dom~{}}f}$ if there is $\boldsymbol{s}\in\mathbb{R}^{n}$ such that for every $\boldsymbol{d}\in\mathbb{R}^{n}$,
$f^{\prime}(\boldsymbol{x},\boldsymbol{d})=\left\langle\boldsymbol{s},\boldsymbol{y}\right\rangle.$
If $\boldsymbol{s}$ exists, then it is unique, it is called the gradient of $f$ at $\boldsymbol{x}$, and it is written as $\boldsymbol{s}\equiv\nabla f(\boldsymbol{x})$.
@@
    The function $f$ has a subgradient $\boldsymbol{s}$ at a point $\boldsymbol{x}\in\mathrm{dom~{}}f\iff f^{\prime}(\boldsymbol{x};\boldsymbol{d})\geqslant\langle\boldsymbol{s},\boldsymbol{d}\rangle$ for every $\boldsymbol{d}\in\mathbb{R}^{n}$ [54, Theorem 23.2, page 216].
@@
    The function $f$ has a unique subgradient at $\boldsymbol{x}$ if and only if $f$ is differentiable at $\boldsymbol{x}$ and $\partial f(\boldsymbol{x})=\{\nabla f(\boldsymbol{x})\}$ [26, Proposition 5.3, page 23].
  • (iv)
@@
If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then its subdifferential is a maximal monotone mapping [55, Theorem 12.17]. If also $f$ is $t$-strongly convex with $t>0$, then for every $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\mathbb{R}^{n}$ and every $\boldsymbol{v}_{1}\in\partial f(\boldsymbol{x}_{1})$, $\boldsymbol{v}_{2}\in\partial f(\boldsymbol{x}_{2})$ we have the stronger monotone inequality [54, Corollary 31.5.2],
@@
Let $f\in\Gamma_{0}(\mathbb{R}^{n})$. The convex conjugate $f^{*}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ of $f$ is defined by
$f^{*}(\boldsymbol{s})\coloneqq\sup_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\langle\boldsymbol{s},\boldsymbol{x}\right\rangle-f(\boldsymbol{x})\right\}.$
By definition, the function $f^{*}$ is in $\Gamma_{0}(\mathbb{R}^{n})$ [26, Definition 4.1, page 17].
@@
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is coercive if for every sequence $\{\boldsymbol{x}_{k}\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ with $\lim_{k\to+\infty}\left\|{\boldsymbol{x}_{k}}\right\|_{2}=+\infty$,
$\lim_{k\to+\infty}f(\boldsymbol{x}_{k})=+\infty.$
If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then $f$ is coercive if and only if $\boldsymbol{0}\in\mathrm{ri}\,\mathrm{dom~{}}f^{*}$ [55, Theorem 11.8, page 479].
@@
Let $C$ be a nonempty, closed and convex subset of $\mathbb{R}^{n}$ and let $\boldsymbol{x}\in\mathbb{R}^{n}$. We denote by $\mathrm{proj}_{C}(\boldsymbol{x})$ the projection of $\boldsymbol{x}$ on $C$:
$\mathrm{proj}_{C}(\boldsymbol{x})\coloneqq\operatorname*{arg\,min}_{\boldsymbol{y}\in C}\left\|{\boldsymbol{x}-\boldsymbol{y}}\right\|_{2}.$
The projection of $\boldsymbol{x}$ on $C$ exists, is unique, and satisfies the characterization

### v2

@@ 6 Numerical experiments
This section presents some numerical experiments to compare the accuracy and run times of Algorithm 1 with some state-of-the-art algorithms for computing regularization and solution paths to (LASSO). Specifically, we compare our algorithms with the glmnet software package glmnet [46, 54, 25, 26], MATLAB’s native lasso implementation mlasso (ostensibly the implementation of the glmnet package to MATLAB, except that it has more options, including for better control of the accuracy), and the fast iterative shrinkage thresholding algorithm fista with a strong selection rule [5, 22, 47, 41]. For the special case $t=0$, we compare the performance of Algorithm 1 with regularization path starting from $(t^{(0)},\boldsymbol{p}^{(0)})=(\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty},-\boldsymbol{b}/t^{(0)})$ and no regularization path starting from input $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. All experiments were carried out in MATLAB on an Apple M3 processor.
@@
All our algorithms require solving a NNLS problem. For this, we use Meyer’s algorithm [36], an active set NNLS algorithm generalizing the Lawson–Hanson algorithm [32] to arbitrary starting active sets, and so particularly suitable for Algorithm 1. For glmnet, we use the code available at https://github.com/junyangq/glmnet-matlab. For fista, we use the variant described in [1, Algorithm 5, Page 197] together with the selection rule described in [47, Equation 7]. For our numerical experiments, we use the benchmark datasets provided by Lorenz et al. [34, Tables 1-2, Section 4], tailored specifically for (LASSO) and (BP) and to use for comparing different numerical algorithms. Each dataset comprises a triplet $(\bm{A},\boldsymbol{b},\boldsymbol{x}_{\text{BP}}^{s})$, where $\boldsymbol{x}_{\text{BP}}^{s}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{m}}\left\|{\boldsymbol{x}}\right\|_{1}\ \text{s.t.}\ \bm{A}\boldsymbol{x}=\boldsymbol{b}$. We use in total ten different datasets, corresponding to some of their largest dense matrix (m=1024,n=8192m=1024,\,n=8192) with six different observed data and solution vectors and their largest sparse matrix (m=8192,n=49152m=8192,\,n=49152) with four different observed data and solution vectors. The solution vectors have either high dynamic range (HDR) or low dynamic range (LDR), and their support either satisfy the so-called Exact Recovery Condition (ERC) [49], an extended form of it (extERC) or, for dense matrices only, no exact recovery condition (noERC) (see [34] for details). All our MATLAB implementations, external software and datasets will be bundled and made publicly available on Zotero.
@@ 6.1 Accuracy checks
We compare how accurately glmnet, fista and mlasso compute solutions to (LASSO) and (dLASSO) (via the optimality conditions $t\boldsymbol{p}^{s}(t,\boldsymbol{b})=\bm{A}\boldsymbol{x}^{s}(t,\boldsymbol{b})-\boldsymbol{b}$) across solution paths. We use dataset 548 (dense $1024\times 8192$ matrix, LDR, noERC) and dataset 474 (sparse $8192\times 49512$ matrix, LDR, extERC) from [34], their most computationally demanding datasets. For each dataset, we first compute the entire solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ using Algorithm 2 and identify all $K$ kinks in $t$ in the solution path (see Remark 5.1). Then we run Algorithm 1, glmnet, fista, and mlasso (for dataset 584 as mlasso does not support sparse matrices) at the kinks $\{t^{(k)}\}_{k=0}^{K}$, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$.
We perform two runs for each dataset. In the first run, we use the default options of glmnet and mlasso and a relative tolerance of $10^{-4}$ for fista (i.e., the algorithm stops when the relative difference in $\ell_{\infty}$-norm of the dual updates is less than $10^{-4}$). In the second run, we use harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista). In both runs, we use a tolerance of $10^{-8}$ in Algorithms 1 and 2 when evaluating their equicorrelation sets on line 3. After convergence of each algorithm, we evaluate the dual objective function in (dLASSO) and number of nonzero components of their primal solutions along the solution path.
@@
At default tolerance, mlasso and fista performed reasonably well while glmnet performed poorly. In particular, glmnet produced a numerical solution with many more nonzero components than all other algorithms. This remained the case even in dataset # 474 with a harsh tolerance. Taken together, we see that Algorithm 1 is the clear winner when it comes to accuracy, with mlasso and fista performing well, and with glmnet performing poorly, even when using harsher tolerances.
Figure 1: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution paths. The dataset used is # 548 from [34]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-4}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
Figure 2: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution path. The dataset used is # 474 from [34]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-4}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
@@ 6.2 Run times comparisons
6.2.1 Computing regularization paths
Table 1 shows the run times of Algorithm 1, glmnet, mlasso, and fista for computing regularization paths. The regularization paths were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$. We also included the point $\{0\}$ for Algorithm 1, which the other algorithms cannot handle. Since we wish to compute accurate regularization paths, we used harsh tolerances for each algorithm, just as in Section 6.1 ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista). Table 1 shows that the overall winner is Algorithm 1; it is significantly faster than fista and mlasso, and slightly faster or comparable to glmnet.
@@ 6.2.1 Computing regularization paths
6.2.2 Computing optimal solutions when $t=0$
Table 2 shows the run times of Algorithm 1 with $t=0$ and using as input $\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, and Algorithm 1 with a regularization path. We used the ten aforementioned benchmark datasets from [34], and the regularization paths for Algorithm 1 were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$ and including $\{0\}$. We did not include MATLAB’s native linear programming solver because we found that it was slow and often failed to converge. We also did not include commercial solvers for (BP)because earlier work by Tendero et al. [45] did similar comparisons with a different version of Algorithm 1 (valid for $t=0$ only) and found it superior in performance and accuracy.
Table 2 shows that all the methods have similar running times. In addition, we’ve verified that all methods correctly computed the solution vectors provided in the benchmark datasets. Algorithm 1 with regularization path is often slower because it used a large number of points.
Datasets Alg. 1 (Reg. path to $t=0$) Alg. 1 (direct at $t=0$)
# 147 (d, HDR, ERC) 0.65 0.019
# 148 (d, HDR, extERC) 0.69 0.041
# 274 (d, HDR, noERC) 1.22 0.72
# 421 (d, LDR, ERC) 0.58 0.020
# 422 (d, LDR, extERC) 0.60 0.035
# 548 (d, LDR, noERC) 1.06 0.58
# 199 (s, HDR, ERC) 0.65 0.028
# 200 (s, HDR, extERC) 1.07 1.23
# 473 (s, LDR, ERC) 0.50 0.028
# 474 (s, LDR, extERC) 0.98 1.29
Table 2: Timings (in seconds) for Algorithm 1 with a regularization path and Algorithm 1 with $t=0$ using $\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$ as input. Total times for 512/1024 logarithmically spaced points and including $\{0\}$ (d: dense/s: sparse), averaged over 5 runs.
@@ 7 Conclusion and future work
In this work, we proved that a minimal selection principle from the theory of differential inclusions (Theorem 3.1) enables one to compute an optimal solution of (dLASSO) from the asymptotic limit of the slow system (12). As the results in Section 3 and 4 show, the slow system can be integrated, yielding the slow solution (31). The slow solution converges in finite time and, at convergence, yields the optimal solution to (dLASSO). From it, one can recover an optimal solution to (LASSO). Taken together, these results yielded Algorithm 1. We also presented, in Section 5, a detailed perturbation analysis of the slow system, including its local and global dependence on the hyperparameter and data. The global continuation of the slow solution provided a rigorous homotopy algorithm for the lasso problem. Our numerical experiments showed that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting its key feature that it neither compromises accuracy nor computational efficiency.
While this work focused on the lasso problem, our results yielded a novel solution method for solving a broad class of projected dynamical systems. We therefore expect that our results will be relevant to applications involving variational inequalities and projected dynamical systems. In addition, we expect that our results can be adapted to compute exact or approximate solutions to a broader class of convex polyhedral-constrained optimization problems. This includes: (i) variations of the lasso problem where the $\ell_{1}$ norm is replaced by a polyhedral norm or the function $\boldsymbol{x}\mapsto\left\|{\bm{M}\boldsymbol{x}}\right\|_{1}$ for some appropriate real matrix $\bm{M}$, (ii) other $\ell_{1}$-regularized problems such as logistic regression, Poisson regression, support vector machines and boosting problems, and (iii) extensions to inequality constraints, i.e., linear and quadratic programmming. These problems will be investigated in future work.
@@ Appendix A Mathematical background
We list here important definitions and technical results from convex and functional analysis used in this work. For comprehensive references, we refer the reader to [42, 43, 3, 21, 29, 30]. All vectors and matrices are denoted in bold typeface. Given an $m\times n$ real matrix $\bm{A}$ and a set of indices $\mathcal{E}\subset\{1,\dots,n\}$, we write $\bm{A}_{\mathcal{E}}$ to denote its $m\times|\mathcal{E}|$ submatrix with columns indexed by $\mathcal{E}$ and we write $\bm{A}^{\top}_{\mathcal{E}}$ to denote its transpose. Similarly, given $\boldsymbol{u}\in\mathbb{R}^{n}$, we write $\boldsymbol{u}_{\mathcal{E}}$ to denote its $|\mathcal{E}|$-dimensional subvector indexed by $\mathcal{E}$.
@@
A nonempty set $C\subset\mathbb{R}^{n}$ is a closed convex polyhedron if it can be expressed as $C\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:\langle\boldsymbol{u}_{j},\boldsymbol{x}\rangle\leqslant r_{j}\ \text{for every $j\in\{1,\dots,l\}$}\right\}$, where $\{\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{l}\}\subset\mathbb{R}^{n}$ and $\{r_{1},\dots,r_{l}\}\subset\mathbb{R}$.
@@
A function $f$ defined over $\mathbb{R}^{n}$ is proper if its domain $\mathrm{dom~}f\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:f(\boldsymbol{x})<+\infty\right\}$ is nonempty and $f(\boldsymbol{x})>$-$\infty$ for every $\boldsymbol{x}\in\mathrm{dom~}f$.
@@
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is lower semicontinuous at $\boldsymbol{x}\in\mathrm{dom~}f$ if for every sequence $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ converging to $\boldsymbol{x}$, $\liminf_{k\to+\infty}f(\boldsymbol{x}_{k})\geqslant f(\boldsymbol{x})$. We say that $f$ is lower semicontinuous if it is lower semicontinuous at every $\boldsymbol{x}\in\mathrm{dom~}f$.
@@
Let $f$ be a proper function with $\mathrm{int}{~\mathrm{dom~}f}\neq\emptyset$. The function $f$ is differentiable at $\boldsymbol{x}\in\mathrm{int}{~\mathrm{dom~}f}$ if there is $\boldsymbol{s}\in\mathbb{R}^{n}$ such that for every $\boldsymbol{d}\in\mathbb{R}^{n}$, $f^{\prime}(\boldsymbol{x},\boldsymbol{d})=\left\langle\boldsymbol{s},\boldsymbol{y}\right\rangle$. If $\boldsymbol{s}$ exists, then it is unique, it is called the gradient of $f$ at $\boldsymbol{x}$, and it is written as $\boldsymbol{s}\equiv\nabla f(\boldsymbol{x})$.
@@
    The function $f$ has a unique subgradient at $\boldsymbol{x}$ if and only if $f$ is differentiable at $\boldsymbol{x}$ and $\partial f(\boldsymbol{x})=\{\nabla f(\boldsymbol{x})\}$ [21, Proposition 5.3, page 23].
@@
If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then its subdifferential is a maximal monotone mapping [43, Theorem 12.17]. If also $f$ is $t$-strongly convex with $t>0$, then for every $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\mathbb{R}^{n}$ and every $\boldsymbol{v}_{1}\in\partial f(\boldsymbol{x}_{1})$, $\boldsymbol{v}_{2}\in\partial f(\boldsymbol{x}_{2})$ we have the stronger monotone inequality [42, Corollary 31.5.2],
@@
Let $f\in\Gamma_{0}(\mathbb{R}^{n})$. The convex conjugate $f^{*}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ of $f$ is defined by $f^{*}(\boldsymbol{s})\coloneqq\sup_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\langle\boldsymbol{s},\boldsymbol{x}\right\rangle-f(\boldsymbol{x})\right\}$. In particular, $f^{*}\in\Gamma_{0}(\mathbb{R}^{n})$ [21, Definition 4.1].
@@
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is coercive if for every sequence $\{\boldsymbol{x}_{k}\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ with $\lim_{k\to+\infty}\left\|{\boldsymbol{x}_{k}}\right\|_{2}=+\infty$, we have $\lim_{k\to+\infty}f(\boldsymbol{x}_{k})=+\infty$. If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then $f$ is coercive if and only if $\boldsymbol{0}\in\mathrm{ri}\,\mathrm{dom~}f^{*}$ [43, Theorem 11.8, page 479].
@@
Let $C$ be a nonempty, closed and convex subset of $\mathbb{R}^{n}$ and let $\boldsymbol{x}\in\mathbb{R}^{n}$. The projection of $\boldsymbol{x}$ on $C$ is $\mathrm{proj}_{C}(\boldsymbol{x})\coloneqq\operatorname*{arg\,min}_{\boldsymbol{y}\in C}\left\|{\boldsymbol{x}-\boldsymbol{y}}\right\|_{2}$. The projection of $\boldsymbol{x}$ on $C$ exists, is unique, and satisfies the characterization